OpenStudy (anonymous):

Which rate is the best deal for a box of cereal? A. $2.49 for 14 ounces B. $3.49 for 18 ounces C. $4.29 for 24 ounces D. $6.39 for 39 ounces

OpenStudy (anonymous):

\[\text{Round}\left[\left\{\frac{2.49}{14},\frac{3.49}{18},\frac{4.29}{24},\frac{6.39}{39}\right\},0.01\right]\to \{0.18,0.19,0.18,0.16\} \]D. 16 cents/oz